Determine the value of the following:



1. P36


2. D5


3. Q3


1
Expert's answer
2022-01-10T10:50:56-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 Class & f_i & Class & cf \\
& & boundaries & & & & \\ \hline
 73-75 & 2 & 72.5-75.5 & 2 \\ \hdashline
 76-78 & 6 & 75.5-78.5 & 8 \\ \hdashline
 79-81 & 12 & 78.5-81.5 & 20 \\ \hdashline
 82-84 & 16 & 81.5-84.5 & 36 \\ \hdashline
 85-87 & 18 & 84.5-87.5 & 54 \\ \hdashline
 88-90 & 39 & 87.5-90.5 & 93 \\ \hdashline
 91-93 & 36 & 90.5-93.5 & 129 \\ \hdashline
 94-96 & 21 & 93.5-96.5 & 150 \\ \hdashline
 97-99 & 5 & 96.5-99.5 & 155 \\ \hdashline
 & n=155 & & & & \\ 
\end{array}"


1.

"P_{36}" class:

Class with "(\dfrac{36n }{100})^{th}" value of the observation in "cf" column

"(\dfrac{36(155) }{100})^{th}" value "=(55.8)^{th}" value

It lies in the class "88-90"

"P_{36}" class: "87.5-90.5"

The lower boundary point of "87.5-90.5" is "87.5."

"L=" ​lower boundary point of median class"=87.5"

"P_{36}=L+\dfrac{\dfrac{36n}{100}-cf}{f}\cdot c=87.5+\dfrac{55.8-54}{39}\cdot 3"

"=87.6385"

2.

"D_5" class:

Class with "(\dfrac{5n }{10})^{th}" value of the observation in "cf" column

"(\dfrac{5(155) }{10})^{th}" value "=(77.5)^{th}" value

It lies in the class "88-90"

"D_5" class: "87.5-90.5"

The lower boundary point of "87.5-90.5" is "87.5."

"L=" ​lower boundary point of median class"=87.5"

"D_5=L+\dfrac{\dfrac{5n}{10}-cf}{f}\cdot c=87.5+\dfrac{77.5-54}{39}\cdot 3"

"=89.3077"



3.

"Q_3" class:

Class with "(\dfrac{3n }{4})^{th}" value of the observation in "cf" column

"(\dfrac{3(155) }{4})^{th}" value "=(166.25)^{th}" value

It lies in the class "91-93"

"Q_3" class: "90.5-93.5"

The lower boundary point of "90.5-93.5" is "90.5."

"L=" ​lower boundary point of median class"=90.5"


"Q_3=L+\dfrac{\dfrac{3n}{4}-cf}{f}\cdot c=90.5+\dfrac{116.25-93}{36}\cdot 3"

"=92.4375"